OpenStudy (anonymous):

Find d2y/dx2 by implicit differentiation, given x^3 + y^3 = 6. Please help!

OpenStudy (amistre64):

how is implicit any different then unimplicit?

OpenStudy (amistre64):

what the derivative of x^3?

OpenStudy (anonymous):

3x^2

OpenStudy (amistre64):

close, but you threw out your derived bit .... bad bad bad 3x^2 x'

OpenStudy (amistre64):

i never said to derive it with respect to x

OpenStudy (amistre64):

what is the derivative of y^3?

OpenStudy (anonymous):

you don't need the x'?

OpenStudy (amistre64):

there is nothing special about an x ....

OpenStudy (amistre64):

what is the derivative of y^3 :)

OpenStudy (anonymous):

and the derivative of y^3 derived implicitly is 3y^2 dy/dx or 3y^2 y'

OpenStudy (amistre64):

right; 3y^2 y'

OpenStudy (amistre64):

soo ... 3x^3 x' + 3y^2 y' = 0 agreed?

OpenStudy (amistre64):

err.. 3x^2 lol

OpenStudy (anonymous):

haha yes

OpenStudy (anonymous):

I know how to get the first derivative, it ends up being -(x^2)/(y^2)

OpenStudy (anonymous):

but for some reason I cannot get the 2nd derivative...

hero (hero):

Amistre, should I just post wolfram's version on here?

OpenStudy (amistre64):

now lets clean it up; with respect to x; x'=1 sooo 3x^2 + 3y^2 y' = 0 ; and solve for y' 3y^2 y' = -3x^2 y' = -3x^2/3y^2 ; and simplify y' = -x^2/y^2 ; and simplify

OpenStudy (amistre64):

if you want, its a free post ;)

hero (hero):

I agree. It's free of charge

OpenStudy (amistre64):

you know the quotient rule right?

OpenStudy (anonymous):

Yes I do... I mean I've tried it and I apparently don't arrive at the right answer, so I'm probably not doing something right. That's why I need help! haha

OpenStudy (amistre64):

lets see where I mess it up at then ;)

hero (hero):

gn, what grade are you in?

OpenStudy (anonymous):

why does that matter? lol

OpenStudy (anonymous):

but I have a meeting to go to that starts at 7! If you guys could help me with this problem real quick I would greatly appreciate it!!!

OpenStudy (amistre64):

\[y' = \frac{-x^2}{y^2}\] \[y' = \left(\frac{-x}{y}\right)^2\] \[y'' = 2\left(\frac{-x}{y}\right)*\frac{d}{dx}\frac{-x}{y}\] \[y'' = 2\left(\frac{-x}{y}\right)*\frac{-x'y+xy'}{y^2}\] so far so good?

hero (hero):

show off

OpenStudy (amistre64):

lol im a genius

OpenStudy (anonymous):

yesss

OpenStudy (amistre64):

\[y'' = 2\left(\frac{-x}{y}\right)*\frac{-x'y+xy'}{y^2}\] \[y'' = \frac{-2x}{y}\frac{(-y+xy')}{y^2}\] \[y'' = \frac{-2x(-y+xy')}{y^3}\] and recall what y' was .... -x^2/y^2 right? \[y'' = \frac{-2x(-y+x\frac{-x^2}{y^2})}{y^3}\]

hero (hero):

I have no idea how he types all of that so fast.

hero (hero):

Some secret program

OpenStudy (anonymous):

this all looks good and makes sense so far

OpenStudy (amistre64):

\[y'' = \frac{2x(y+x\frac{x^2}{y^2})}{y^3}\] \[y'' = \frac{2x(y^3+x^3)}{y^5}\] \[y'' = \frac{2xy^3+2x^4}{y^5}\] and i think that as simple as it goes

OpenStudy (amistre64):

whatd you come up with?

OpenStudy (anonymous):

I came up with that except I I don't get how you go from that first step to the second one... so how does (y+x (x^2/y^2)) become (y^3+x^3)? Probably a dumb question but I don't see how that works

OpenStudy (amistre64):

x*x^2 = x^3 right?

OpenStudy (amistre64):

then to add fraction you need like denominators like they teach you back in kindergarten; so y^2 it up

OpenStudy (amistre64):

y*y^2 = y^3

OpenStudy (amistre64):

the denom of it sinks under to merge into the y^3

OpenStudy (anonymous):

I just put that answer in and it was wrong :(

OpenStudy (amistre64):

..... then it prolly aint as simplified as it wants

hero (hero):

amistre's getting rusty

OpenStudy (amistre64):

rusty? i mess up all the time, this is par for the course :)

OpenStudy (amistre64):

wolfram aint computing my input right

OpenStudy (anonymous):

So for a similar problem: Find d2y/dx2 by implicit differentiation, given x^3 + y^3 = 8. the answer is -16x/(y^5)

OpenStudy (amistre64):

yeah, looks like i got the y^5 part

OpenStudy (amistre64):

the first still goes to -x^2/y^2

OpenStudy (amistre64):

\[y'=\frac{-x^2}{y^2}\] \[y''=\frac{y^2(-x^2)'-(y^2)'(-x^2)}{y^4}\] \[y''=\frac{y^2(-2x)-(2y\ y')(-x^2)}{y^4}\] \[y''=\frac{-2xy^2+(2x^2y\ y')}{y^4}\]

OpenStudy (amistre64):

\[y''=\frac{-2xy^2+(2x^2y\ \frac{-x^2}{y^2})}{y^4}\] \[y''=\frac{-2xy^2+(2x^2\ \frac{-x^2}{y})}{y^4}\] \[y''=\frac{-2xy^2-(2x^2\ \frac{x^2}{y})}{y^4}\] \[y''=\frac{-2xy^2-(2x^4\ \frac{1}{y})}{y^4}\] \[y''=\frac{-2xy^2*\frac{y}{y}-(2x^4\ \frac{1}{y})}{y^4}\] \[y''=\frac{-2xy^3\frac{1}{y}-2x^4\ \frac{1}{y}}{y^4}\] \[y''=\frac{-2xy^3-2x^4}{y^5}\] hmmmm

OpenStudy (amistre64):

what does y^3=?

OpenStudy (anonymous):

Sorry... I am really late for my meeting now... I can't stay on here any longer. Thank you so much for your help

OpenStudy (amistre64):

x^3 + y^3 = 8 y^3 = 8-x^3 ..right?

OpenStudy (amistre64):

\[y''=\frac{-2x(8-x^3)-2x^4}{y^5}\] \[y''=\frac{-16x+2x^4-2x^4}{y^5}\] \[y''=\frac{-16x}{y^5}\]

OpenStudy (amistre64):

that did it

OpenStudy (amistre64):

it wasnt that it was wrong, it just wasnt as simplified as they wanted it to be